OpenStudy (anonymous):

7r+22k=101 8r+20k=100

OpenStudy (anonymous):

What are you trying to do with the equations? Combine them?

OpenStudy (anonymous):

or finding the values of r and k?

OpenStudy (whpalmer4):

Assuming you're trying to find the values of \(r, k\) that satisfy the system of equations: \[7r+22k=101\]\[8r+20k=100\] I would suggest multiplying everything on the first line by -8, and everything on the second line by 7. After you do so, you get: \[-56r - 176k = -808\]\[56r + 140k = 700\]Adding down each column, you are left with a simpler equation in terms of \(k\) which you should be able to solve. After doing so, plug the value you found for \(k\) into one of the original equations and solve for \(r\).